OpenStudy (anonymous):

4(sin x)^2 + 2 (cos x)^2 = 3. How do I solve the equation in the interval [0, 2pi) ?

OpenStudy (anonymous):

you can replace \[\cos^2(x)\] by \[1-\sin^2(x)\] to get an equation only in sine

OpenStudy (anonymous):

re-write as: \[2\sin ^{2}x +2\cos ^{2}x + 2\sin ^{2}x=3\] \[2(\sin ^{2}x+\cos ^{2}x)+2\sin ^{2}x=3\] can you pick up from that?

OpenStudy (anonymous):

\[2+2\sin ^{2}x=3\] \[2\sin ^{2}x = 1\] finish it :)

OpenStudy (anonymous):

\[sinx=+/- 1/\sqrt{2}\] so, x=+/- pi/4

OpenStudy (anonymous):

how did you get rid of the cosine?

OpenStudy (anonymous):

sin^2 + cos^2 = 1

OpenStudy (anonymous):

ohh okay i see now

OpenStudy (anonymous):

thank youuuu :)

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

Please click: 'Good Answer' button on the right... it will remove this problem from unanswered :)